Calculate the predicted wavelength of a photon that is produced due to the transition of an electron from the 5th shell to the fourth shell.
Answer #1

Here photon will be emitted

1/lambda = R* (1/nf^2 - 1/ni^2)

R is Rydberg constant. R = 1.097*10^7

1/lambda = R* (1/nf^2 - 1/ni^2)

1/lambda = 1.097*10^7* (1/4^2 - 1/5^2)

lambda = 4.051*10^-6 m

lambda = 4051 nm

Answer: 4.05*10^3 nm